# algorithmic

### v1

@@ 6.2 Experiment Results
General Performance Comparison: From Figure 3 (a) and (b), which compare the performance across different method families under policy shift, we can see that the DM-PS family consistently outperforms other methods substantially. In Table 4 in Appendix D.2, we further present the number of conditions where our proposed method families perform optimally. The DM-PS family emerges as the top performer in both known and unknown logging policies, outperforming other baseline methods in over 90% of the conditions. For general covariate shift settings, when the density ratio $\mathcal{W}(\bm{x},a)$ is known (Figure 3 (e)), the DM-GCS family exhibits the best performance. When $\mathcal{W}(\bm{x},a)$ is unknown (Figure 3 (f)), the DM-PS family leads in performance. In both scenarios (Figure 3 (e) and (f)), both DM-GCS and DM-PS families outperform SnIPS-GCS in over 70% of the examined conditions.
Comparison between DM-PS and DM-GCS: When comparing DM-PS and DM-GCS in Figure 3 (g) and (h) as well as in Table 2, we observe that DM-PS and its family perform better when the density ratio $\mathcal{W}(\bm{x},a)$ is unknown, while DM-GCS and its family excels when $\mathcal{W}(\bm{x},a)$ is known. These results can be attributed to that the GCS setting introduces the disparity between $P_{s}(\bm{x})$ and $P_{t}(\bm{x})$, creates larger distribution shifts and harder-to-estimate density ratios, and destabilizes the training of DM-GCS and the IPS component within DR methods. We further demonstrate this by comparing SnIPS and SnIPS-GCS in Table 5 in Appendix D.3. Also, the relatively weaker performance of DM-GCS when $\mathcal{W}(\bm{x},a)$ is unknown can be attributed to the estimation of $P_{s}(\bm{x},a)$, which requires a concurrent estimation of both $P_{s}(\bm{x})$ and $\beta(a|\bm{x})$. The inherent difficulty in accurately estimating $P_{s}(\bm{x})$, which is typically more complex than estimating the logging policy, negatively affects the performance of the DM-GCS family.
@@
Policy Shift. Consider two types of softened shift logging policies: small shifts, such as $\pi_{(0.95,0)}(a|\bm{x})$ and $\pi_{(0.7,0.1)}(a|\bm{x})$, where logging policy deviations from the target policy are modest; and large shifts, represented by $\pi_{(0.5,0.1)}(a|\bm{x})$ and $\pi_{(0.1,0)}(a|\bm{x})$, that exhibit more substantial deviations. The Tweak-1 and Dirichlet policies are treated as even larger shift logging policies due to more probability values close to 0 or 1. Figure 4 (a) illustrates the proportion of each DM method achieving the best performance across these conditions, revealing that DM-PS consistently surpasses other methods across diverse logging policy scenarios. DM-PS’s advantage grows as the shift enlarges, demonstrating its excellence in accommodating policy shifts in reward estimation and aligning well with our theoretical results in Section 5.1.1.
General Covariate Shift. Figure 4 (b) showcases the CDF of the relative MSE of DM-GCS with respect to DM-PS in Gaussian covariate shift scenarios, with a larger $\omega$ indicating a larger covariate shift. As the covariate shift increases, DM-GCS exhibits widening advantages over DM-PS, indicating the benefits of accommodating covariate shifts in reward estimation.
@@ 7 CONCLUSION AND FUTURE WORK
Appendix A Description of More Baseline Methods
@@
Appendix B Derivation of Robust Regression Model
@@
Appendix C Proof of Bias Analysis
@@ C.1 Consistency
Appendix D Experimental details
We here describe the logging policy, target policy, and covariate we mentioned in Section 6.1 in detail and the specific parameters we used. We also provide experiment results on each dataset. The code repository for the experiments can be found at https://anonymous.4open.science/r/Distributionally-Robust-Policy-Evaluation-under-General-Co–variate-Shift-in-Contextual-Bandits-6532/README.md
D.1 Data Generation
@@
D.2 Comparison of DM-PS Family and Other Baselines under Policy Shift
@@
D.3 Comparison of SnIPS and SnIPS-GCS under different covariate shift
@@
D.4 Experiments Results on Each Dataset

### v2

@@ 6.2 Experiment Results
General Performance Comparison: From Figure 3 (a) and (b), which compare the performance across different method families under policy shift, we can see that the DM-PS family consistently outperforms other methods substantially. In Table 4 in Appendix E.2, we further present the number of conditions where our proposed method families perform optimally. The DM-PS family emerges as the top performer in both known and unknown logging policies, outperforming other baseline methods in over 90% of the conditions. For general covariate shift settings, when the density ratio $\mathcal{W}(\bm{x},a)$ is known (Figure 3 (e)), the DM-GCS family exhibits the best performance. When $\mathcal{W}(\bm{x},a)$ is unknown (Figure 3 (f)), the DM-PS family leads in performance. In both scenarios (Figure 3 (e) and (f)), both DM-GCS and DM-PS families outperform SnIPS-GCS in over 70% of the examined conditions. In (c), (d), (g), and (h), DM-PS and DM-GCS significantly outperform MRDR and DRPE. Further comparisons with advanced OPE methods such as SWITCH and DRoS are provided in Appendix E.4.
Comparison between DM-PS and DM-GCS: When comparing DM-PS and DM-GCS in Figure 3 (g) and (h) as well as in Table 2, we observe that DM-PS and its family perform better when the density ratio $\mathcal{W}(\bm{x},a)$ is unknown, while DM-GCS and its family excels when $\mathcal{W}(\bm{x},a)$ is known. These results can be attributed to that the GCS setting introduces the disparity between $P_{s}(\bm{x})$ and $P_{t}(\bm{x})$, creates larger distribution shifts and harder-to-estimate density ratios, and destabilizes the training of DM-GCS and the IPS component within DR methods. We further demonstrate this by comparing SnIPS and SnIPS-GCS in Table 5 in Appendix E.3. Also, the relatively weaker performance of DM-GCS when $\mathcal{W}(\bm{x},a)$ is unknown can be attributed to the estimation of $P_{s}(\bm{x},a)$, which requires a concurrent estimation of both $P_{s}(\bm{x})$ and $\beta(a|\bm{x})$. The inherent difficulty in accurately estimating $P_{s}(\bm{x})$, which is typically more complex than estimating the logging policy, negatively affects the performance of the DM-GCS family.
@@
1. Policy Shift. Consider two types of softened shift logging policies: small shifts, such as $\pi_{(0.95,0)}(a|\bm{x})$ and $\pi_{(0.7,0.1)}(a|\bm{x})$, where logging policy deviations from the target policy are modest; and large shifts, represented by $\pi_{(0.5,0.1)}(a|\bm{x})$ and $\pi_{(0.1,0)}(a|\bm{x})$, that exhibit more substantial deviations. The Tweak-1 and Dirichlet policies are treated as even larger shift logging policies due to more probability values close to 0 or 1. Figure 4 (a) illustrates the proportion of each DM method achieving the best performance across these conditions, revealing that DM-PS consistently surpasses other methods across diverse logging policy scenarios. DM-PS’s advantage grows as the shift enlarges, demonstrating its excellence in accommodating policy shifts in reward estimation and aligning well with our theoretical results in Section 5.1.1.
2. General Covariate Shift. Figure 4 (b) showcases the CDF of the relative MSE of DM-GCS with respect to DM-PS in Gaussian covariate shift scenarios, with a larger $\omega$ indicating a larger covariate shift. As the covariate shift increases, DM-GCS exhibits widening advantages over DM-PS, indicating the benefits of accommodating covariate shifts in reward estimation.
@@ 7 CONCLUSION AND FUTURE WORK
Acknowledgement
This work is partially supported by the JHU-Amazon AI2AI faculty award, the Discovery Award of the Johns Hopkins University, and a seed grant from JHU Institute of Assured Autonomy.
Appendix A Related Work
DR methods (Dudik et al., 2011; 2014; Wang et al., 2017; Farajtabar et al., 2018) strike a balance between the biased, low-variance DM and the unbiased, high-variance IPS by incorporating both. It can also be interpreted as using control variates within an IPS framework (Thomas & Brunskill, 2016; Veness et al., 2011). If either IPS or DM is unbiased, DR methods are guaranteed to be unbiased (Dudik et al., 2011). Much of the follow-up work on DR methods has focused on mitigating the detrimental effects of variance or extreme probabilities from the IPS component. For example, SnDR utilizes SnIPS (Swaminathan & Joachims, 2015b), SWITCH estimator truncates IPS using a threshold (Wang et al., 2017), and DR with Optimistic Shrinkage (DRoS) finds a weighting strategy by optimizing a sharp bound on mean squared error (Su et al., 2019a). While much of the prior work has emphasized improving the IPS component of DR estimators and their asymptotical guarantees, our research targets enhancing the robustness of conditional reward estimators, especially under large shifts and limited data. Our key insight is to view this problem as a covariate shift problem and utilize robust regression to train the conditional reward estimator. Our proposed method can also be naturally incorporated into existing methods like DMs and DRs by offering a distributionally robust reward estimator.
Distributionally robust optimization (DRO) (Ben-Tal et al., 2009; Słowik & Bottou, 2022), a prevalent approach to handling distribution shifts, introduces the idea of optimizing with respect to the most adverse scenarios over a defined family of distributions. This has led to different variants such as Wasserstein DRO (Blanchet et al., 2021). A few studies have explored distributional robustness in policy evaluation. Si et al. (2020) and Kallus et al. (2022) augmented evaluation robustness by conservatively estimating reward mapping with KL-divergence uncertainty sets. However, in these methods, the uncertainty set covers all possible shifts in the joint distribution of context, action, and reward, which does not characterize the covariate shift specifically. As a result, the estimator can be overly robust and less effective. Another line of work utilize DRO for estimating confidence interval in the off-policy setting (Karampatziakis et al., 2020; Dai et al., 2020; Faury et al., 2020), following a generalized empirical likelihood approach (Duchi et al., 2021). The differences from our approach lie in the formulation of the DRO problem. Specifically, they do not use the conditional distribution of the target variable (the reward in our case) as the adversarial player, and the uncertainty set constraint is formulated by the perturbation of empirical distribution of the data with distance measured using f-divergence, rather than using feature matching as in our case. Robust regression (Chen et al., 2016; Liu & Ziebart, 2017) tailored distributional robustness methods specifically for the setting of covariate shift. It constructs a predictor that approximately matches training data statistics but is otherwise the most uncertain on the testing distribution by minimizing the worst-case expected target loss and obtaining a parametric form of the predicted output labels’ probability distributions. We are the first to employ robust regression methods for policy evaluation, enabling the refinement of direct methods and the further enhancement of doubly robust estimators.
Appendix B Description of More Baseline Methods
@@
Appendix C Derivation of Robust Regression Model
@@
C.1 Gaussian Predictive Form
Here we provide an example that the data is generated from cubic form distribution instead of Gaussian distribution. In this scenario, we can adjust the robust regression model by incorporating the cubic function feature constraint instead of the quadratic constraint. Denote $x$ as the covariate variable, $y$ as the target variable, $f_{0}(y|x)$ as the prior and $\phi(x,y)$ to be the cubic function feature constraint. We can obtain the derived predictive model form of $p(y|x)\propto f_{0}(y|x)\exp(-w(x)\theta\phi(x,y))$, where $\theta$ is the model parameter and $w(x)$ is the density ratio. This estimator has a hard-to-track normalization denominator for the predictive model distribution, so we need to sample from this distribution to obtain the prediction of $y$ from such a model, which will sacrifice the nice implementation convenience of our robust regression model, and we may need to tailor specific sampling methods to efficiently obtain the prediction from such an estimator.
We conduct experiments under settings where $p(y|x)$ is not Gaussian by generating synthetic data $(x,y)$ such that $p(y|x)\propto\exp(y^{3}-2y^{2}x-3yx^{2}-4y^{2}+5yx+6y)$ following a cubic feature constraint assumption. Specifically, we created the distribution shift by generating $x$ following $x_{\text{train}}\sim N(0,1)$ and $x_{\text{test}}\sim N(1,2)$. There are 500 data points in both the train and test datasets. Since the cubic feature constraint does not result in a Gaussian distribution predictor model like the quadratic feature setting and the denominator of such distribution is intractable, we use discrete target $y$ by sampling $y\in$0,1} from the aforementioned prediction distribution in this experiment, for both the data generation and getting predictions from the robust regression model.
The cubic form robust regression achieved a mean squared error (MSE) of $0.31(\pm 0.00029)$. The quadratic form robust regression received an MSE of $0.42(\pm 0.0003)$, and the linear regression model received an MSE of $0.49(\pm 1.2\times 10^{-10})$. These results are averaged over 10 runs. These results suggest that a more tailored feature set could improve the performance in synthetic settings and using a simpler hypothesis class could induce bias. However, given the original model for regression without Gaussian assumption, $p(y|x)\propto\exp(-w(x)\theta\phi(x,y))$, can be less convenient for making predictions (which is the reason we need to use sampling to get the $y$ in this synthetic experiments) and the broad applicability of Gaussian data assumptions, a robust regression model with Gaussian assumptions is still preferred in practice.
Appendix D Proof of Bias Analysis
@@ D.1 Consistency
Appendix E Experimental details
We here describe the logging policy, target policy, and covariate we mentioned in Section 6.1 in detail and the specific parameters we used. We also provide experiment results on each dataset. The code for the experiments can be found at https://github.com/guoyihonggyh/Distributionally-Robust-Policy-Evaluation-under-General-Co–variate-Shift-in-Contextual-Bandits
E.1 Data Generation
@@
E.2 Comparison of DM-PS Family and Other Baselines under Policy Shift
@@
E.3 Comparison of SnIPS and SnIPS-GCS under different covariate shift
@@
E.4 Further Comparison with Advanced OPE Methods
We present the result of our method applied to and compared with other advanced OPE approaches such as SWITCH (Wang et al., 2017) and DRoS (Su et al., 2020) in Figure 5. Our method can be integrated into SWITCH and DRoS by replacing their reward estimator with our estimators, DM-PS and DM-GCS. Incorporating DM-PS into SWITCH and DRoS results in SWITCH-PS and DRoS-PS, respectively. Similarly, applying DM-GCS to SWITCH and DRoS results in SWITCH-GCS and DRoS-GCS. As shown in Figure 5, SWITCH-PS and DRoS-PS consistently outperform the original SWITCH and DRoS methods, respectively, across all cases, including both policy shift and general covariate shift cases. Additionally, in the general covariate shift setting, the performance of DRoS is also further improved by DM-GCS. These results demonstrates the advantage of integrating DM-PS and DM-GCS into SWITCH and DRoS.
Figure 5: Subfigures (a) and (b) present the experiment results under policy shift. Subfigures (c) and (d) present the experiment results under general covariate shift. DM-PS and DM-GCS enhance the performance of SWITCH and DRoS.
E.5 Experiments Results on Each Dataset